# Temporal Action Localization for Inertial-based Human Activity Recognition

### v1

@@ 3.3. Architectures Overview
In order to predict activity segments $Y=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$ within an input video, the ActionFormer, TemporalMaxer and TriDet model all follow the same sequence labeling problem formulation for action localization. That is, given a set of feature input vectors $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}$, a model aims to classify each timestamp’s as either one of the activity categories $C$ or as background (NULL-class) and regress the distance towards the timestamp’s corresponding segment’s start and end point. More specifically, given the input vectors $X$ a model aims to learn to label the sequence as
(12) $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}\to\mathbf{\hat{Y}}=\{\mathbf{\hat{y}_{1}},\mathbf{\hat{y}_{2}},...,\mathbf{\hat{y}_{T}}\}$
@@
, where $\mathbf{\hat{y}_{t}}=(p(a_{t}),d_{t}^{s},d_{t}^{e})$ at timestamp $t$ is defined by a probability vector $p(a_{t})$ indicating the class-wise probability of the timestamp being classified as one of the activity categories $C$ and $d_{t}^{s}>0$ and $d_{t}^{e}>0$ being the distance between the current timestamp $t$ and the current segment’s onset and offset. Note that $d_{t}^{s}$ and $d_{t}^{e}$ are not defined if the timestamp is to be classified as background. The sequence labeling formulation can then be easily decoded to activity segments as:
@@
ActionFormer
Introduced in 2022, the ActionFormer was among the first single-stage, anchor-free TAL models which followed a transformer-based design (Zhang et al., 2022b). The architecture, depicted in Figure 3, follows an encoder-decoder structure that decomposes the previously defined task of learning to predict activity segments using a function $f$ as $h\circ g$, where $g:X\to Z$ serves as the encoder and $h:Z\to\hat{Y}$ is the decoder. As the encoder Zhang et al. (2022b) employ a combination of convolutional layers and a transformer network. The convolutional layers, paired with ReLU activation functions, project each input feature vector $\mathbf{x}_{t}$ into a $D$-dimensional space. To limit complexity, each layer of the transformer network applies alternating layers of multi-head local self-attention and MLP. To capture information at different temporal scales, the transformer network is structured in such a way that across its $L$ layers, the initially embedded feature vector $\mathbf{z}_{0}$ at timestamp $t$ is downsampled using a strided depthwise 1D-convolution.
After encoding the input sequence into a feature pyramid $Z=\{\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{L}\}$, which captures information at various temporal scales, the decoder decodes the feature pyramid $Z$ into sequence labels $\hat{Y}$. To do so, the decoder includes both a classification and regression head. The classification head, being 1D convolutional network complemented with layer normalization and ReLU activation, predicts for all $L$ pyramid levels the probability vector $p(a_{t})$ at each timestep $t$. The regression head, following the same design, predicts the onset ($d_{t}^{s}$) and offset distance ($d_{t}^{e}$) at all pyramid levels $L$ for each timestamp $t$. Note that for both the classification and regression heads, parameters are shared across pyramid levels and output regression ranges are pre-specified per pyramid level. Finally, to reduce overlapping instances and create a final set of predicted activity segments, the obtained predictions across all pyramid levels are postprocessed using Soft-NMS (Bodla et al., 2017).
Figure 3. Architecture overview of the ActionFormer proposed by Zhang et al. (2022b). The architecture follows a encoder-decoder structure. The encoder, comprising of a projection using convolutional layers and a transformer network, encodes input sequences into a feature pyramid, which captures information at various temporal scales. The decoder, consisting of a classification and regression head, then decodes each timestamp within the feature pyramid to sequence labels, i.e. a class probability vector and the timestamp’s activity onset and offset distance. Note that weights of the decoder are shared across pyramid levels.
TemporalMaxer
Building upon the architectural design of the Actionformer (Zhang et al., 2022b), the TemporalMaxer proposed by Tang et al. (2023) suggests modifying the encoder of the ActionFormer to employ solely max pooling and remove any multi-head attention layers (see Figure 4). According to the authors, the removal of the transformer-based layers does not come at the cost of a lost in information, as the extracted feature embeddings $X$ outputted by the projection layer are already discriminative enough to achieve state-of-the-art predictive performance. Same as the ActionFormer, the weights of both the classification and regression heads within the TemporalMaxer are shared across pyramid levels, output regression ranges are pre-specified per pyramid level, and to create a final set of predicted activity segments, obtained predictions across all pyramid levels are postprocessed using Soft-NMS (Bodla et al., 2017).
Figure 4. Architecture overview of the TemporalMaxer (Tang et al., 2023). The authors propose to modify the transformer of the ActionFormer (Zhang et al., 2022b) to employ solely max pooling, removing any self-attention mechanisms from the architecture. Each resulting embedded timestamp within the feature pyramid is then decoded using a classification and a regression head. Note that weights of the decoder are shared across pyramid levels.
TriDet
Similar to the TemporalMaxer, the TriDet model (see Figure 5) suggested to make modifications to the architectural design of the ActionFormer by altering both the encoder and regression head decoder (Shi et al., 2023). The authors claim that a transformer-based encoder is prone to suffer from rank loss while simultaneously being high in computational complexity. Consequently, Shi et al. (2023) propose replacing the projection and transformer network of the ActionFormer with Scalable-Granularity Perception (SGP) layers. These layers, being fully-convolutional, comprise of two branches: an instant-level branch, which increases the discriminability between the activity classes and the NULL-class, and a window-level branch, aiming to capture semantic context from a wider receptive field. Mathematically, the SGP operator replacing the mulit-head attention operator is defined as:
(14) $f_{SGP}=\phi(x)FC(x)+\psi(x)(Conv_{w}(x)+Conv_{kw}(x))+x$
, where $FC$ are fully-connected, $Conv_{w}$ are 1D-depth-wise convolutional layers over the window size $w$ and $k$ is a scalable factor controlling the granularity of captured temporal information. In the TriDet model, the $L$ layers of the feature pyramid $Z=\{\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{L}\}$ are thus defined as first downsampling the input feature embeddings $L$ times and applying a SGP layer to each downsampled embedding. According to the authors, the ActionFormer architecture suffers from imprecise boundary predictions. To address this, the original regression head is replaced by a Trident-head, which aims to overcome the ActionFormer’s shortcomings via an estimated relative probability distribution around each activity boundary. The Trident-head consists of three main components: a start head, responsible for locating the start of an activity segment; an end head, responsible locating the end of an activity segment; and a center-offset head, responsible for locating the center of an activity segment. Using these three heads, the Trident-head predicts the start ($d_{t}^{s}$) and end boundary ($d_{t}^{e}$) for all $L$ pyramid levels for each timestamp $t$. Note that all heads, including the classification head, are modelled as three 1D-convolutional layers with parameters shared across all feature pyramid levels.
Figure 5. Architecture overview of the TriDet model proposed by Shi et al. (2023). Compared to the Actionformer (Zhang et al., 2022b), the architecture proposes the following changes: 1. replace the projection and transformer network of the ActionFormer with Scalable-Granularity Perception (SGP) layers along with preceding downsample operators via max-pooling, and 2. replace the regression head with a Trident head, consisting of a start, end and center-offset head. Each resulting embedded timestamp within the feature pyramid is then decoded using the newly defined decoder. Note that weights of the decoder are shared across pyramid levels.
3.4. Similarities and Differences between Inertial and TAL Models
@@
Similar to inertial models, the ActionFormer, TemporalMaxer and TriDet model aim to learn temporal dependencies by learning discriminative feature embeddings. Nevertheless, do the authors accredit much of their models’ performance to the constructed multi-layer feature pyramid. As described above, the feature pyramid within each of the three models downsamples the input embedding multiple times in order to create embedding vectors of different temporal granularity. Though this is similar to downsampling performed by concatenating multiple convolution layers without padding, as demonstrated by all inertial architectures in this benchmark analysis, the TAL models use each embedding vector separately instead of only using the resulting embedding vector of the last pyramid layer. More specifically, the authors claim that predicting timestamps at different degrees of temporal granularity holds significant value and is essential in order to correctly predict different types and lengths of activities. Furthermore, as already detailed in previous chapters, the TAL community aims to localize segments rather than provide window-wise predictions. Due to their objective thus differing from that of the inertial community, TAL models are trained not only on a classification loss, but also a regression loss, which optimizes each timestamps corresponding activity onset and offset. Lastly, the use of self-attention based layers has shown to improve results in both communities, yet variations of the ActionFormer show that these layers are not necessarily needed. More specifically, the use of fully-convolutional SGP layers in the TriDet model show that, rather than the type of technique being important, focus needs to be put specifically learning both local and global temporal information for each timestamp.
@@ 4.1. Datasets
We evaluate each algorithm featured in this benchmark analysis using 6 popular HAR datasets, namely the Opportunity (Roggen et al., 2010), SBHAR (Reyes-Ortiz et al., 2016), Wetlab (Scholl et al., 2015), WEAR (Bock et al., 2023), Hang-Time (Hoelzemann et al., 2023) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016). The datasets, all covering different application and recording scenarios, provide us with a challenging set of prediction problems to properly assess the strengths and weaknesses of each model. Table 1 summarizes the key characteristics of each dataset. In addition to vital information such as participant count, activity count, and sensor axes count, the table also includes details on the type of activities found in each dataset. We classify activities into four types: (1) periodic activities, characterized by recurring periodic patterns; (2) non-periodical activities consisting of non-occurring, non-periodical patterns; and (3) complex activities, defined by an arbitrary sequence of non-periodic and periodic activities. Lastly, following the works of Alwassel et al. (2018), we provide average count of segments across all subjects, categorizing each segment within each dataset into five bins: XS: (0 seconds, 3 seconds], S: (3 seconds, 6 seconds], M: (6 seconds, 12 seconds], L: (12 seconds, 18 seconds], and XL: more than 18 seconds.
Table 1. Key characteristics of the datasets used in this benchmark analysis. The table provides: participant count (#Sbjs), activity count (#Cls), sensor axes count (#Axes), average number of segments per participant based on absolute length of the segment (#Segments) and type of activities found in the dataset.
Dataset #Sbjs #Cls #Axes #Segments (Alwassel et al., 2018) Type of Activities
@@
Opportunity (Roggen et al., 2010) 4 17(+1) 113 $403\pm 63$ $185\pm 48$ $48\pm 10$ $3\pm 2$ $0\pm 0$ non-periodic, complex
SBHAR (Reyes-Ortiz et al., 2016) 30 12(+1) 3 $3\pm 2$ $9\pm 2$ $6\pm 4$ $10\pm 4$ $13\pm 3$ periodic, non-periodic
Wetlab (Scholl et al., 2015) 22 8(+1) 3 $7\pm 5$ $7\pm 3$ $7\pm 2$ $5\pm 2$ $13\pm 2$ complex
WEAR (Bock et al., 2023) 18 18(+1) 12 $1\pm 1$ $1\pm 1$ $1\pm 1$ $1\pm 1$ $30\pm 9$ periodic
Hang-Time (Hoelzemann et al., 2023) 24 5(+1) 3 $154\pm 53$ $22\pm 7$ $10\pm 7$ $3\pm 3$ $3\pm 2$ non-periodic, periodic, complex
RWHAR (Sztyler and Stuckenschmidt, 2016) 15 8 21 $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $9\pm 1$ periodic
Opportunity
Introduced by Roggen et al. (2010) in 2010, the Opportunity dataset features a sensor-rich recording scenario of four participants performing various activities of daily living. As part of this benchmark analysis we chose to use only the sensor subset associated with the Opportunity challenge Chavarriaga et al. (2013), which consists of 113 sensor axes from IMU and accelerometer body-worn sensors, all sampling at 30Hz. The Opportunity challenge provides two prediction scenarios, i.e. types of annotations, that is modes of locomotion and gestures. For the benchmark analysis presented in this paper, we focus on the latter, which involves predicting 17 types of activities. The Opportunity dataset was recorded under laboratory-like conditions having each participant take part in six different recording sessions. There were two types of sessions: ADL sessions, repeated 5 times, in which participants followed a higher-level protocol and were allowed to interleave activities, and Drill sessions, repeated once, in which participants performed a sequence of activities 20 times. Breaks were included in all recording sessions, resulting in a substantial NULL-class being part of the dataset.
SBHAR
Comprising of 30 participants performing six activities of daily living, the SBHAR dataset augments the base locomotion activities with annotated transitional activities, representing transition periods between two base activities (Reyes-Ortiz et al., 2016). The dataset itself consists of uninterrupted 3D-accelerometer data sampled at 50 Hz, which was captured by a smartphone attached to the waist of each participant. Similar as the Opportunity dataset, the SBHAR dataset was also recorded under laboratory-like conditions, with participants following an experimental protocol that defined repetition count, order and execution length of each activity.
Wetlab
Published by Scholl et al. (2015), the Wetlab dataset provides 3D-accelerometer data from 22 participants. A single inertial sensor, placed in a fixed orientation on the dominant wrist of each participant, captured data at a constant sampling rate of 50Hz. The recording of the dataset took place in a wetlab environment in which each participant performed two DNA extraction experiments. Participants were allowed to take as much time as needed for each experiment. However, following an experimental wetlab protocol, recordings of different participants contain almost identical sequences of consecutive activities. Nevertheless, since not all steps in the experimental protocol were mandatory, some activities were skipped by certain participants. The provided annotated activity streams are two-fold, with one set indicating the steps within the recording protocol and the other representing the recurring base activities, such as peeling, pestling, and cutting. During our benchmark analysis the latter will be used, being the 8 base activities (+ the NULL-class), as it is most comparable to the prediction scenarios present in the other benchmark datasets used.
WEAR
Published by Bock et al. (2023) in 2023, WEAR consists of 3D-accelerometer and egocentric video data from 18 participants performing 18 workout sports activities at 10 outdoor locations. These activities include running-, stretching- and strength-based exercises, with base activities like push-ups being complemented with complex variations that alter and/ or extend the performed movement during the exercise. The 18 activities were divided across multiple recording sessions, with each session consisting of uninterrupted data streams of all modalities, i.e. including a NULL-class. Although participants were assigned a fixed set of workout activities and instructed to perform each activity for approximately 90 seconds, each participant had the freedom to choose the order of activities and take breaks as desired. During each workout session, four inertial sensors all sampling at 50Hz were placed on both left and right wrists and ankles.
Hang-Time
Published by Hoelzemann et al. (2023), Hang-Time features uninterrupted data streams of 24 participants during a basketball practice. The 3D-accelerometer data of each participant was sampled using a smartwatch worn on each participant’s right wrist. Although the dataset provides both locomotion and basketball-related activity labels across three different types of sessions (drills, warm-up and game), we intentionally exclude the locomotion activities during our benchmark analysis, replacing them with a NULL-class label. In the drill sessions, participants followed a strict experimental protocol with predefined activity durations and sequences. In contrast, the warm-up and game sessions allowed participants to commence activities at their discretion in a more real-world recording scenario.
RWHAR
Consisting of 15 participants performing a set of 8 different locomotion activities, the RWHAR dataset (Sztyler and Stuckenschmidt, 2016) was recorded at in-the-wild locations. Unlike the other datasets featured in this benchmark analysis, the RWHAR dataset does not contain a NULL-class, as recordings were trimmed to only contain the activities of relevancy, cutting out the breaks in-between. Except for jumping, each activity was performed the same amount of time, i.e. around 15 minutes. Participants wore a total of 6 body-worn sensors, all set to sample at 50Hz. During this benchmark analysis, we will only be using the 3D-accelerometer and 3D-gyroscope data provided to make modalities comparable to that of the other datasets.
@@ 4.2. Training Pipeline
All experiments were initiated employing a Leave-One-Subject-Out cross-validation. This type of validation involves iteratively training on all but one participant’s data and using the hold-out participant’s data during validation until all participants have been evaluated, ensuring that each network architecture is assessed based on its capabilities to generalize across unseen participants. All datasets were windowed using a sliding window of one second with a 50% overlap across windows. For all inertial architectures (Ordóñez and Roggen, 2016; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022) we employed a similar optimization as proposed with the release of the shallow DeepConvLSTM, namely an Adam optimizer paired with a learning rate of $1e^{-4}$, a weight decay of $1e^{-6}$, and Glorot weight initialization (Glorot and Bengio, 2010). To allow each model to converge more properly, we increase the number of epochs to 100 and employ a step-wise learning-rate schedule that multiplies the learning rate by a factor of 0.9 after every 10 epochs. For all architectures, we fixed the hidden dimension of the recurrent layers to employ 128 units and scaled kernel sizes of the convolutional filters according to the relative difference in sampling rate among the different input datasets. In line with how the Attend-and-Discriminate architecture was first introduced, we optimized said architecture using center-loss (Wen et al., 2016) as opposed to a weighted cross-entropy loss, which was used during training of all other inertial architectures. Lastly, as proposed by the authors, we do not shuffle batches during the training of the shallow DeepConvLSTM.
@@
For all three TAL architectures (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) part of this benchmark analysis, we chose to employ hyperparameters reported by the authors that produced best results on the EPIC-KITCHENS dataset (Damen et al., 2022). The hyperparamenters, though optimized for a different modality than inertial data (egocentric videos), have shown to produce competitive results on the WEAR dataset (Bock et al., 2023) and are thus considered a good starting point for evaluating the applicability of the three architectures on other HAR datasets. Nevertheless, given the small size of the tested inertial datasets compared to datasets used by the TAL community, we chose to increase the amount of epochs to 100 throughout all TAL-based experiments. As TAL models are designed to predict both a regression (boundary prediction) and classification target (segment label), loss calculation of the ActionFormer, TemporalMaxer and TriDet models are performed as a weighted combination of a regression loss (IoU loss (Rezatofighi et al., 2019)) and classification loss (focal loss (Li et al., 2020)). The latter, originates from the object detection community, which, like HAR, often times suffers from the background class dominating the class distribution. Furthermore, as we treat each subject’s data as a continuous data stream, we apply the TAL models to predict each subject’s data as a whole. With the input sequences thus being of arbitrary length, the ActionFormer, TemporalMaxer and TriDet model limit their complexity by randomly truncating the input sequence to not exceed a maximum sequence length during training. Furthermore, center sampling is employed when training the three TAL models. This technique makes it such that only time steps within an interval around a segment’s center are to be considered positive, while all other parts are considered negative. Within the TAL community, center sampling has shown to encourage higher scores around activity centers and overall improve results (Tian et al., 2022; Zhang et al., 2020).
@@
As inertial models are tasked to predict on a per-window basis and, in most cases, are trained using shuffled data, the architectures suffer from producing only small coherent segments. Caused by frequently occuring activity switches, the fragmentation ultimately leads to significantly lower mAP scores being produced by the inertial architectures as opposed to the TAL models. Therefore, to remove only short lasting switches, predictions of inertial-based architectures mentioned in this paper were smoothed using a majority-vote filter. The exact size of the majority filter was chosen dataset-specific, being, for example, 15 seconds in the case of the WEAR dataset. Similar to the inertial models, the TAL architectures are tasked to predict class probabilities and segment boundaries of each windowed timestamp. Consequently, without applying any confidence threshold, all predicted activity segments are considered during creation of the prediction timeline causing accuracy of the NULL-class to be significantly low. Therefore, to alleviate this, we apply an optimized confidence threshold on predicted segments of all TAL models. Similarly to the majority filter, the score threshold for each architecture was chosen dataset-specific, being, for instance, 0.2 in the case of the WEAR dataset. This eliminates low scoring segments and substantially lowers the confusion of the NULL-class with the other activities.

### v2

@@ 3.1. TAL Architectures Overview
In order to predict activity segments $Y=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$ within an input video, the ActionFormer, TemporalMaxer and TriDet model all follow the same sequence labeling problem formulation for action localization. That is, given a set of feature input vectors $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}$, a model aims to classify each timestamp as either one of the activity categories $C$ or as background (or null) class and regress the distance towards the timestamp’s corresponding segment’s start and end point. More specifically, given the input vectors $X$ a model aims to learn to label the sequence as
(2) $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}\to\mathbf{\hat{Y}}=\{\mathbf{\hat{y}_{1}},\mathbf{\hat{y}_{2}},...,\mathbf{\hat{y}_{T}}\},$
@@
where $\mathbf{\hat{y}_{t}}=(p(a_{t}),d_{t}^{s},d_{t}^{e})$ at timestamp $t$ is defined by a probability vector $p(a_{t})$ indicating the class-wise probability of the timestamp being classified as one of the activity categories $C$ and $d_{t}^{s}>0$ and $d_{t}^{e}>0$ being the distance between the current timestamp $t$ and the current segment’s onset and offset. Note that $d_{t}^{s}$ and $d_{t}^{e}$ are not defined if the timestamp is to be classified as background. The sequence labeling formulation can then be easily decoded to activity segments with:
@@
The authors of the ActionFormer, TemporalMaxer, and TriDet models attribute much of their models’ performance to the constructed multi-layer feature pyramid. The feature pyramid within each of the three models downsamples the input sequence of sliding windows multiple times to create representations of a participant’s data stream at different temporal granularities (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023). Using this method, the TAL models can learn both short and long temporal dependencies across a participants’s timeline, as the temporal distance between two embedding vectors within the feature pyramid increases the further we move down the pyramid. This use of downsampling to capture different length temporal patterns is comparable to how inertial models concatenate convolutional layers without padding to increase the receptive field of each convolutional kernel. However, both types of models differ in the types of patterns learned. The feature pyramid of the TAL models enables them to learn cross-window, temporal patterns of arbitrary length, while the convolutional parts of the inertial models enable them to learn in-window temporal patterns of arbitrary length.
@@
Due to the TAL community’s training objective differing from that of the inertial community, TAL models are trained not only on a classification loss, but also a regression loss, which optimizes each timestamp’s corresponding activity onset and offset. The use of self-attention based layers has shown to improve results in both communities, yet variations of the ActionFormer show that these layers are not necessarily needed. More specifically, the use of fully-convolutional SGP layers in the TriDet model show that, rather than the type of technique being important, focus needs to be put specifically on learning both local and global temporal information for each timestamp.
Figure 4. Architecture overview of the ActionFormer proposed by Zhang et al. (2022b). The architecture follows a encoder-decoder structure. The encoder encodes input sequences into a feature pyramid, which captures information at various temporal scales. The decoder, consisting of a classification and regression head, then decodes each timestamp within the feature pyramid to sequence labels, i.e. a class probability vector and the timestamp’s activity onset and offset distance. The TriDet(Shi et al., 2023) and TemporalMaxer(Tang et al., 2023) both follow the same encoder-decoder structure, yet suggest architectural changes.
[…]
Visualization of the ActionFormer architecture. Inputs, i.e. 1-dimensional feature embeddings each representing a sliding window of a subject’s data stream, are passed through an encoder which consists of projection and transformer layers. Each layer of the feature pyramid of the encoder is passed to a decoder which predicts class probabilities, onset and offset of each data point within the feature pyramid. Weights of the decoder are shared across the feature pyramid layers.
@@ 4.1. Datasets
We evaluate each algorithm featured in this benchmark analysis using 6 popular HAR datasets, namely the Opportunity (Roggen et al., 2010), SBHAR (Reyes-Ortiz et al., 2016), Wetlab (Scholl et al., 2015), WEAR (Bock et al., 2023), Hang-Time (Hoelzemann et al., 2023) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016). The datasets, all covering different application and recording scenarios, provide us with a challenging set of prediction problems to properly assess the strengths and weaknesses of each model. Table 1 summarizes the key characteristics of each dataset. In addition to vital information such as participant count, activity count, sampling rate and sensor axes count, the table also includes details on the scenario and type of activities found in each dataset. We classify activities into four types: (1) periodic activities, characterized by recurring periodic patterns; (2) non-periodical activities consisting of non-occurring, non-periodical patterns; and (3) complex activities, defined by an arbitrary sequence of non-periodic and periodic activities. Lastly, following the works of Alwassel et al. (2018), we provide average count of segments across all subjects, categorizing each segment within each dataset into five bins: XS: (0 seconds, 3 seconds], S: (3 seconds, 6 seconds], M: (6 seconds, 12 seconds], L: (12 seconds, 18 seconds], and XL: more than 18 seconds.
Table 1. Key characteristics of the datasets used in this benchmark analysis. The table provides: participant count (#Sbjs), activity count (#Cls), sampling rate (SR), sensor axes count (#Axes), average number of segments per participant based on absolute length of the segment (#Segments), overall scenario (e.g. activities of daily living (ADL)) and type of activities found in the dataset.
Dataset #Sbjs #Cls SR #Axes #Segments (Alwassel et al., 2018) Scenario Type of Activities
@@
Opportunity (Roggen et al., 2010) 4 17(+1) 30 113 $403\pm 63$ $185\pm 48$ $48\pm 10$ $3\pm 2$ $0\pm 0$ ADL non-periodic, complex
SBHAR (Reyes-Ortiz et al., 2016) 30 12(+1) 50 3 $3\pm 2$ $9\pm 2$ $6\pm 4$ $10\pm 4$ $13\pm 3$ Locomotion (non-)periodic
Wetlab (Scholl et al., 2015) 22 8(+1) 50 3 $7\pm 5$ $7\pm 3$ $7\pm 2$ $5\pm 2$ $13\pm 2$ Laboratory complex
WEAR (Bock et al., 2023) 18 18(+1) 50 12 $1\pm 1$ $1\pm 1$ $1\pm 1$ $1\pm 1$ $30\pm 9$ Sports periodic
Hang-Time (Hoelzemann et al., 2023) 24 5(+1) 50 3 $154\pm 53$ $22\pm 7$ $10\pm 7$ $3\pm 3$ $3\pm 2$ Sports (non-)periodic, complex
RWHAR (Sztyler and Stuckenschmidt, 2016) 15 8 21 50 $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $9\pm 1$ Locomotion periodic
@@ 4.2. Training Pipeline
Prediction Scenarios
All experiments were initiated employing a Leave-One-Subject-Out cross-validation. This type of validation involves iteratively training on all but one participant’s data and using the hold-out participant’s data during validation until all participants have been evaluated, ensuring that each network architecture is assessed based on its capabilities to generalize across unseen participants. All datasets were windowed using a sliding window of one second with a 50% overlap across windows. In order to allow for fair comparison, our benchmark analysis assesses the TAL models employing two prediction scenarios, with the differences in their predicted output between inertial-based and TAL models. Being the scenario TAL models were intended to be used for, within the first prediction scenario (Offline Activity Recognition) the TAL models are tasked to predict each participant’s data as a whole, i.e. one batch representing the full data stream available of one participant. Nevertheless, to also assess the models’ capabilities to provide window-level predictions, within the second prediction scenario (Online Activity Recognition) we artificially chunk participants’ timelines making the models predict the timelines in chunks (e.g., 30 seconds). As inertial-based models unlike TAL models are designed to make window-based predictions, the second prediction scenario will assess the TAL models online capabilities while also benchmarking them with removed temporal context. Note that in case of the Offline Activity Recognition we test the TAL models using both a single-stage as well as two-stage training (as described in Chapter 3).
Hyperparameters
@@
For all inertial architectures (Ordóñez and Roggen, 2016; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022) we employed a similar optimization as proposed with the release of the shallow DeepConvLSTM, namely an Adam optimizer paired with a learning rate of $1e^{-4}$, a weight decay of $1e^{-6}$, and Glorot weight initialization (Glorot and Bengio, 2010). To allow each model to converge more properly, we increase the number of epochs to 100 and employ a step-wise learning-rate schedule that multiplies the learning rate by a factor of 0.9 after every 10 epochs. For all architectures, we fixed the hidden dimension of the recurrent layers to employ 128 units and scaled kernel sizes of the convolutional filters according to the relative difference in sampling rate among the different input datasets. In line with how the Attend-and-Discriminate architecture was first introduced, we optimized said architecture using center-loss (Wen et al., 2016) as opposed to a weighted cross-entropy loss, which was used during training of all other inertial architectures. Lastly, as proposed by the authors, we do not shuffle batches during the training of the shallow DeepConvLSTM.
@@
The three TAL architectures (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023), we chose to employ hyperparameters reported by the authors that produced best results on the EPIC-Kitchens dataset (Damen et al., 2022). The hyperparamenters, though optimized for a different modality than inertial data (egocentric videos), have shown to produce competitive results on the WEAR dataset (Bock et al., 2023) and are thus considered a good starting point for evaluating the applicability of the three architectures on other HAR datasets. Nevertheless, given the small size of the tested inertial datasets compared to datasets used by the TAL community, we chose to increase the amount of epochs to 100 throughout all TAL-based experiments. As TAL models are designed to predict both a regression (boundary prediction) and classification target (segment label), loss calculation of the ActionFormer, TemporalMaxer and TriDet models are performed as a weighted combination of a regression loss (IoU loss (Rezatofighi et al., 2019)) and classification loss (focal loss (Li et al., 2020)).
Postprocessing
To compare predictions of both TAL and inertial-based models, segments (TAL) and windowed-predictions (inertial-based) first needed to be translated back to sample-wise predictions. In case of the segments we do so by translating each segment ”overwrite” the prediction timeline with its associated label. By starting with the segment associated with the lowest prediction confidence, segments with a high prediction confidence are preferred in case of overlapping segments. In case of the inertial-based architectures, we start by iterating over the windowed predictions in order of occurrence in the timeline having them determine the samples they are associated with. To deal with the overlap amongst windows, preceding windows are allowed to ”overwrite” the label of samples of previous windows which they are overlapping with.
@@
As inertial models are tasked to predict on a per-window basis and, the architectures suffer from frequently occurring activity switches ultimately leading to fragmented segments which significantly lower mAP scores being produced as opposed to the TAL models. Therefore, to remove only short lasting switches, predictions of inertial-based architectures mentioned in this paper were smoothed using majority-vote filters. The exact size of the majority filter was chosen dataset-specific, determined via trying out a selection of filters between 2.5 and 40 seconds. Specifically, the filters were chosen as: 2.5 seconds (Opportunity), 20 seconds (Wetlab), 5 seconds (SBHAR), 15 seconds (WEAR), 5 seconds (Hang-Time) and 40 seconds (RWHAR). Similar to the inertial models, the TAL architectures are tasked to predict class probabilities and segment boundaries of each windowed timestamp. Consequently, without applying any confidence threshold, all predicted activity segments are considered during creation of the prediction timeline causing accuracy of the NULL-class to be significantly low. Therefore, to alleviate this, we apply an optimized confidence threshold on predicted segments of all TAL models. Similarly to the majority filter, the score threshold for each architecture was chosen dataset-specific, determined via trying out a selection of thresholds between 0.05 and 0.5 seconds. This eliminates low scoring segments and substantially lowers the confusion of the NULL-class with the other activities. More details on the effect of the majority filter as well as score thresholding can be found in the supplementary material.